HUBBLE SPACE TELESCOPE DAILY REPORT #2111 PERIOD COVERED: 0000Z (UTC) 04/27/98 - 0000Z (UTC) 04/28/98 Daily Status Report as of 118/0000Z 1.0 ACTIVITIES SCHEDULED AND ACCOMPLISHED: 1.1 Completed WF/PC-2 6776 (A Comparative HST Imaging Study of Radio Galaxies and Hosts) A clear understanding of the nature of the host galaxies of radio-quiet quasars, radio-loud quasars and radio galaxies is of central importance in any attempt to unify and relate the three main classes of powerful active galaxies. Existing HST studies of quasar hosts, whilst producing many interesting new results, have yielded surprisingly little hard information on the properties of the hosts themselves. This imaging study og radio galaxy 3C436 using the WF/PC-2 will remedy this situation by undertaking redder, line-free observations of our three matched AGN samples. The proposal completed with no reported problems. 1.2 Completed Three Sets of NIC/2/3 7907 (NICMOS Pure Parallel Winter 97-98) The Near Infrared Camera (Cameras Two and Three) was used to make pure parallel spectrographic observations of sky regions when another instrument is prime. The observations were completed as planned, and no anomalies were reported. 1.3 Completed Two Sets of STIS/CCD 7908 (POMS Test Proposal: STIS Non-Scripted Parallel Proposal) The Space Telescope Imaging Spectrograph (CCD) was used to make calibration observations as part of the POMS Test Proposal. This was a STIS non-scripted parallel proposal. The observations were completed as planned, and no problems were reported. 1.4 Completed Three Sets of NIC/2 7226 (Search for Massive Jupiters) The Near Infrared Camera (Camera Two) was used to make a corographic search for massive planets around nearby, young main sequence star HD177996. Because of the extreme youth of these objects, any low-mass brown dwarf and planetary companions will still be in a higher luminosity phase and thus easily detectable. A dark calibration image was also taken. The observations were executed as scheduled, and no anomalies were noted. 1.5 Completed FGS 7493 (Calibrating the Mass-Luminosity Relation at the End of the Main Sequence) The FGSs were used to make observations of star G250-29 in order to make calibration observations of the mass-luninocity relation at the end of the main sequence. The study utilizes HST-FGS and infrared speckle imaging to calibrate the mass-luminosity relation for stars less massive than 0.2 Msun, with special emphasis on objects near the stellar/brown dwarf border. The observations were completed as scheduled, and no problems were noted. 1.6 Completed Three Sets of WF/PC-2 7712 (WF/PC-2 Cycle 7 Supplemental Darks pt2/3) The WF/PC-2 was used to make a series of Cycle 7 supplemental calibration darks. This dark calibration program obtains three dark frames every day to provide data for monitoring and characterizing the evolution of hot pixels. The observations were executed as scheduled, and no anomalies were noted. 1.7 Completed STIS/CCD 7930 (STIS CCD Binned Bias Monitor) The Space Telescope Imaging Spectrograph (CCD) was used to make numerous bias exposures in 1 x 2 binned mode in order to permit construction of "superbias" frames in which the effects of read noise have been rendered negligible. The observations were executed as scheduled, and no anomalies were noted. 1.8 Completed STIS/MA2 7647 (STIS NUV-MAMA Cycle 7 Flats) The Space Telescope Imaging Spectrograph (MA2) was used to make near-ultraviolet (NUV) flat calibration images. This program will obtain NUV-MAMA flat-field observations with the D2 lamp for the construction of on-orbit D-flats for G230M. A dark calibration image was also taken. The observations were executed as scheduled, and no anomalies were noted. 1.9 Completed Three Sets of WF/PC-2 7620 (WF/PC-2 Cycle 7 Standard Darks) The WF/PC-2 was used to make a series of Cycle 7 standard dark calibration images. This dark calibration program obtains dark frames every week in order to provide data for the ongoing calibration of the CCD dark current rate, and to monitor and characterize the evolution of hot pixels. The observations were executed as scheduled, and no anomalies were noted. 1.10 Completed STIS/CCD 7932 (Spectral Purity and slit throughputs for the First Order Spectroscopic Modes) The Space Telescope Imaging Spectrograph (CCD) was used to make observations of standard star GD71 in order to study the spectral purity and slit throughputs for the first order spectroscopic modes. The purpose of this proposal is to measure and quantify this effect by observing strong absorption lines through slits of different sizes. A wave length calibration image was also taken. The observations were completed as scheduled, and no problems were reported. 1.11 Completed STIS/MA1 7645 (STIS FUV-MAMA Cycle 7 Flats) The Space Telescope Imaging Spaectrograph (MA1) was used to make far-ultraviolet flat calibration images. This program will obtain FUV-MAMA flat-field observations with the Kr lamp for the construction of on-orbit D-flats for G140M. A dark calibration image was also taken. The observations were completed as planned, and no anomalies were reported. 1.12 Completed Two Sets of WF/PC-2 7501 (Expansion Parallax Distances to Planetary Nebulae) The WF/PC-2 was used to make parallax distance measurements of the planetary nebulae IC4663 and NGC-6326. The largest impediment to our understanding of planetary nebulae is the lack of reliable distances. Acquired by generalized statistical methods in all but a handful of cases, individual PNe distances are often uncertain by factors of two or more. As detailed in 2.1, the acquisition for the first iteration of this proposal failed and the take data flag remained down. Nine observations were affected. Otherwise, the observations were executed as scheduled, and no further anomalies were noted. 1.13 Completed STIS/CCD 7803 (Daily Darks to Update Acquisition Bad Pixel Table Part III) The Space Telescope Imaging Spectrograph (CCD) was used to make daily dark calibration updates. These daily darks are intended to identify hot pixels for update to the acquisition on-board FSW hot pixel table. Daily hot pixel lists can also be used for science data analysis. As detailed in 2.1, the acquisition for the first iteration of this proposal failed and the take data flag remained down. Two observations were affected. Otherwise, the observations were completed as planned, and no further anomalies were reported. 1.14 Completed STIS/CCD 7926 (CCD Dark and Bias Monitor -- Continued) The Space Telescope Imaging Spectrograph (CCD) was used to make several dark and bias calibration observations. This is to measure CCD dark current and bias rates for the two supported amplifier settings {gain = 1 and gain = 4} in order to gather data for making superdarks and superbiases. These observations will also track the growth of hot pixels on the CCD in general and within the various target acquisition apertures specifically. The observations were executed as scheduled, and no anomalies were noted. 1.15 Completed STIS/CCD 7911 (STIS Galactic Plane Pure Parallel Program) The Space Telescope Imaging Spectrograph (CCD) was used to make low galactic latitude survey observations in parallel with the prime science observations. Short exposures with the 50CCD imaging mode will provide the wavelength zero-point calibration. These data will be placed immediately into the Hubble Data Archive. The observations were executed as scheduled, and no anomalies were noted. 1.16 Completed NIC/1/2/3 7832 (The Age of Inner Bulge Globular Clusters) The Near Infrared Camera (Cameras One, Two and Three) was used to make observations of the globular cluster Terzan-4 which is located in the Milky Way's central bulge. Until recently very little was known about the age the globular clusters of the galactic bulge, as a combination of crowding, differential reddening, and distance made turnoff photometry beyond the reach of ground based observations. The observations were completed as planned, and no problems were reported. 1.17 Completed WF/PC-2 7282 (A Continuing Snapshot Survey for Compact Stellar Groups and Clusters around Wolf-Rayet Stars) The WF/PC-2 was used to make observations of the Wolf-Rayet star WR104 as part of a continuing snapshot survey of compact stellar groups and clusters. The newly imaged clusters allow us to accurately determine the distances, luminosities, and ages of the WR members. The observations were executed as scheduled, and no anomalies were noted. 1.18 Completed NIC/2 7960 (NICMOS Persistence Test) The Near Infrared Camera (Camera Two) was used to conduct a persistence test. The goal of these tests is to determine whether multiple reads hasten the reduction of image persistence from infrared photons and cosmic rays. The observations were completed as planned, and no anomalies were reported. 1.19 Completed NIC/2/3 7453 (The SBF Hubble Diagram) The Near Infrared Camera (Camera Two and Three) was used to make deep NICMOS observations of galactic cluster Abell-3581 to determine its distance using the surface brightness fluctuation {SBF} method. HST is essential for its high spatial resolution which makes it possible to detect and remove globular clusters and background galaxies that contaminate the SBF measurement. The observations were executed as scheduled, and no anomalies were noted. 1.20 Completed WF/PC-2 7625 (WF/PC-2 Cycle 7 Earth Flats) The WF/PC-2 was used to make a series of Earth flat calibration images. These flat fields will allow mapping of the OTA illumination pattern and will be used in conjunction with previous internal and external flats to generate new pipeline superflats. The observations were completed as planned, and no anomalies were reported. 2.0 FLIGHT OPERATIONS SUMMARY: 2.1 Guide Star Acquisitions: Scheduled Acquisitions : 11 Successful : 10 The acquisition scheduled for 117/1138Z failed to the RGA mode when the scan step limit was exceeded. HSTAR 6450 was written. The two full maneuver updates prior to the acquisition had both failed, causing excessive attitude error (HSTAR 6449). The proposals described in 1.12 and 1.13 were affected. Subsequently, an ARU/PRT was uplinked at 117/123100Z, and the following acquisitions were successful. Scheduled Reacquisitions : 5 Successful : 5 2.2 FHST Updates: Scheduled : 14 Successful : 14 2.3 Operations Notes: Per ROP SR-1, the SSR correctable EDAC counts were cleared three times. At 118/004641Z STIS experienced an MCE reset on MAMA1 during a SAA crossing and while the high voltage was off. HSTAR 6451 was written. The STIS error counter was cleared at 118/010256Z per ROP NS-12. Recovery was completed at 118/015003Z with the high voltage on command in the SMS. Science was not impacted. 3.0 SIGNIFICANT FORTHCOMING EVENTS: Continuation of normal science observations and calibrations. /DMH